A simple pendulum has a period of 3.0seconds. If the value of (g) is taken as 9.9ms-2. Calculate the length of the pendulum


1
Expert's answer
2022-05-04T12:40:24-0400
"T=2\pi\sqrt{\frac{l}{g}}"

"l=\frac{gT^2}{4\pi^2}=\frac{9.9*3.0^2}{4*3.14^2}=2.3\:\rm m"